OpenStudy (anonymous):

The volume of a rectangular box is equal to the length times the width times the height. Write an expression for the volume of a box with length 15 cm, width w cm, and height 10 cm. Be sure to simplify the expression.

OpenStudy (anonymous):

v = 15 * w * 10 v = 150 * w v = 150w